Bellingham Company produces a product that requires six standard pounds per unit. The standard price is $7 per pound. If 3,100 units used 19,200 pounds, which were purchased at $6.79 per pound, what is the direct materials (a) price variance, (b) quantity variance, and (c) cost variance? Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.
Answer #1

Materials Price Variance = (SP-AP) *AQ

= (7 - 6.79)*19,200

= - 4032 Favourable (Minus sign here)

Materials Quantity Variance = (SQ-AQ) *SP

= (3100*6 - 19,200)*7

= 4200 Unfavorable

Cost variance = 4032 F + 4200 U

= 168 Unfavorable